3x + 7y = 1
 x + 2y = 1

A) solution = (4,9)   B) solution = (5,-2) 
C) solution = (1,1/2) D) solution = (1/3 , 1/7)

Answer by AnlytcPhil(1116)

3x + 7y = 1
 x + 2y = 1

A) solution = (4,9)   B) solution = (5,-2) 
C) solution = (1,1/2) D) solution = (1/3 , 1/7) 

===================================================================

3x + 7y = 1
 x + 2y = 1

Make the y's cancel out by multiplying the first 
equation by -2 and the second equation through by +7

-2[3x + 7y = 1]
 7[ x + 2y = 1]

  -6x - 14y = -2
   7x + 14y =  7

Now draw a line and add vertically term by term

  -6x - 14y = -2
   7x + 14y =  7
  覧覧覧覧覧覧覧
    x       =  5

Now make the x's cancel out by multiplying the 
first equation by 1 and the second equation 
through by -3

 1[3x + 7y = 1]
-3[ x + 2y = 1]

   3x + 7y =  1
  -3x - 6y = -3

Now draw a line and add vertically term by term

   3x + 7y =  1
  -3x - 6y = -3
  覧覧覧覧覧覧覧
         y = -2

So (x,y) = (5,-2) and the answer is choice B)

Edwin
AnlytcPhil@aol.com